OpenStudy (anonymous):

Simplify. (4√10 + 3)^2

OpenStudy (anonymous):

Simplify: (y – 4)/(y2 – 16) A. y + 4 B. y – 4 C. y D. 1/(y+4)

OpenStudy (anonymous):

(4√10 + 3)^2 =24√10 +169

OpenStudy (anonymous):

(y – 4)/(y2 – 16)=1/(y+4)

OpenStudy (anonymous):

(4√10 + 3)^2 This is a problem of the form (a + b)^2 = a^ + 2ab + b^2 a =4√10 b = 3 a^2 = 4 * 4 *√10 *√10 =16*10 = 160 b^2 = 9 2ab = 2*3*4√10 = 24√10 Therefore answer = ... 169 + 24√10

OpenStudy (anonymous):

Errata Second line should say (a + b)^2 = a^2 + 2ab + b^2